OpenStudy (morganchavez):

PLEASE HELP DOES ANYONE KNOW HOW TO DO THIS?!!! ALGEBRA 2, UNIT 2, LESSON 3: Properties of Rational Exponents ALTERNATE ASSIGNMENT 2.) Simplify the radical in exponent form. Please show all work.

OpenStudy (morganchavez):

\[\frac{ 1 }{\sqrt[3]{x^-6?} ? }\]

OpenStudy (amistre64):

\[\large\sqrt[n]{k^m}=k^{m/n}\]

OpenStudy (morganchavez):

ok thx